ula is twelve and her grandfather is sixty. In how many years will her grandfather be three times as old as she will be at that time?
Answer by ikleyn(52932)

.
60+x = 3*(12+x)

60 + x = 36 + 3x

60-36 = 3x - x


2x = 24  ====>  x = 24/2 = 12.


Answer.  In 12 years.

Solved.